Calculate the number of pounds of CO2 released into the atmosphere when a 19.0-gallon tank of gasoline is burned in an automobile engine. Assume that gasoline is primarily octane, C8H18, and that the density of gasoline is 0.692 g·mL–1 (this assumption ignores additives). Also assume complete combustion.
Answer #1

The combustion reaction taking place is:

C8H18 + 12.5O2 ---> 8CO2 + 9H2O

Mass of octane taken = Density*Volume = 0.692*71922.8 = 49770.58 g

Moles of octane = Mass/MW = 49770.58/114.23 = 435.71

According to reaction stoichiometry,

Moles of CO2 produced = 8*Moles of octane = 435.71*8 = 3485.68

Mass of CO2 produced = Moles*MW = 3485.68*44 = 153369.92 g = 338.12 pounds

Hope this helps !